# Extremely Amenable Automorphism Groups of Countable Structures

### v1

@@ 7. Concluding Remarks and Further Questions
It is clear that for $\Delta,\Lambda\in\mathsf{CDVS}$, we have $\Delta\approx\Lambda$ iff $\mathcal{A}_{\Delta}\cong\mathcal{A}_{\Lambda}$. This gives a Borel reduction from $\approx_{\mathsf{CDVS}}$ to an isomorphism relation. ∎
It follows that $\approx_{\mathsf{CDVS}}$ is Borel reducible to an equivalence relation induced by a Borel action of $S_{\infty}$. However, we do not know if $\approx_{\mathsf{CDVS}}$ or $\approx$ is Borel.
The main question left open is, of course, about the exact complexity of $\cong_{\mathsf{EA}}$ in the Borel reducibility hierarchy. Here we mention another class of automorphism groups known to be extremely amenable. Recall that an ultrametric is a metric satisfying $d(x,z)\leq\max\{d(x,y),d(y,z)\}$ for any $x,y,z$. An ordering $<$ on an ultrametric space $X$ is convex if for any $x\in X$ and $s>0$, the ball $B(x,s)=\{y\in X:d(x,y)\leq s\}$ is a convex set in the ordering $<$, i.e. if $y_{1}<z<y_{2}$ and $y_{1},y_{2}\in B(x,s)$, then $z\in B(x,s)$. An ultrametric space with a convex ordering is called a convexly-ordered ultrametric space.
@@
Denoting the Fraïssé limit of this class by $B_{S}^{c<}$, by the KPT correspondence, ${\rm Aut}(B_{S}^{c<})$ is extremely amenable. We note that if $S$ and $T$ are countable sets of positive real numbers that are order-isomorphic, then ${\rm Aut}(B_{S}^{c<})$ and ${\rm Aut}(B_{T}^{c<})$ are isomorphic as topological groups. If the converse is true, then we would obtain a new lower bound for the complexity of $\cong_{\mathsf{EA}}$, namely, the most complex one isomorphism relation for any class of countable structures. In particular, $\cong_{\mathsf{EA}}$ as a set would be complete analytic.
Finally, we note that there is a noticeable difference between Theorems 1.2 and 1.5, namely, in Theorem 1.2 we have the equivalence between the isomorphism of the topological groups and the isomorphism of them as abstract groups, while in Theorem 1.5 we do not have this equivalence. The reason for this equivalence to hold in Theorem 1.2 is that $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ has ample generics (see Corollary 4.1 of Solecki [30]), i.e., $G$ has a comeager orbit in the conjugacy action of $G$ on $G^{n}$, for each $n\geq 1$. By a result of Kechris–Rosendal [16], if a Polish group has ample generics, then every abstract homomorphism into a separable group is automatically continuous and thus it has a unique Polish topology. However, for the ordered case, Slutsky [29] showed that ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ does not have ample generics. Hence the question remains whether $\Delta$ and $\Lambda$ are equivalent if $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ and $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Lambda})$ are isomorphic as abstract groups.

### v2

@@ 6. Isomorphism of Extremely Amenable Groups in the Framework of Descriptive Set Theory
Theorem 6.1.
The class $\mathsf{EA}$ of extremely amenable closed subgroups of $S_{\infty}$ is Borel. Consequently, $\cong_{\mathsf{EA}}$ is an analytic equivalence relation.
Proof.
$g\cdot(a_{1},\dots,a_{n})=(g(a_{1}),\dots,g(a_{n})).$
We note that the collection of all $\mathcal{L}$-structures with the universe $\mathbb{N}$ forms a Polish space $X_{\mathcal{L}}=\prod_{n,i\geq 1}X_{n,i}$, where each $X_{n,i}$ is a copy of the Cantor space $2^{\mathbb{N}^{n}}$. The map $G\mapsto\mathcal{A}_{G}$ from $\mathcal{F}(S_{\infty})$ to $X_{\mathcal{L}}$ is Borel. In fact, a subbasic open set in $X_{\mathcal{L}}$ is determined by a tuple $\bar{a}=(a_{1},\dots,a_{n})$ being assigned value 0 or value $1$. If $U$ is a basic open set in $X_{\mathcal{L}}$, then whether $\mathcal{A}_{G}\in U$ is determined by finitely many requirements of form $\exists g\in G\ g(\bar{a})=\bar{b}$ and its negation, where $\bar{a},\bar{b}$ are tuples of the same length. Thus $\{G\in\mathcal{F}(S_{\infty})\colon\mathcal{A}_{G}\in U\}$ is clearly a Borel set.
Let $\mathcal{K}$ be the age of $\mathcal{A}_{G}$. Then $\mathcal{A}_{G}=\mbox{Flim}\,\mathcal{K}$. By Theorem 2.2, $G=\mbox{\rm Aut}(\mathcal{A}_{G})=\mbox{\rm Aut}(\mbox{Flim}\,\mathcal{K})$ is extremely amenable if and only if $\mathcal{K}$ has the Ramsey property and every element of $\mathcal{K}$ is rigid. Thus to prove $\mathsf{EA}$ is Borel, it suffices to show both the following:
  1. (i)
  2. (ii)
After the completion of an earlier version of this paper, we learned from Aleksander Iwanow that Theorem 6.1 had been proved by Iwanow–Majcher-Iwanow in response to a question asked by Nies in the Logic Blog in 2020. The proof is included in the manuscript [24] and is different from the one above.
The following theorem is a corollary of our results about countable distance value sets in the preceding section (in particular, Corollary 5.3).
Theorem 6.2.
$E^{\infty}_{S_{\infty}}$ is Borel reducible to $\cong_{\mathsf{EA}}$. In particular, $\cong_{\mathsf{EA}}$ is not Borel.
Proof.
[…]
  1. (i)
  2. (ii)
  3. (iii)
    $^{\mathcal{A}}$ is a linear order.
Now, given an enumeration of $\mathcal{K}^{<}_{\Delta}$, one can construct the Fraïssé limit in a Borel way. This follows from the existence proof of Fraïssé’s theorem (see, e.g. Lemma 7.1.6 of Hodges [15]). Since the Fraïssé structure is ultrahomogeneous, its automorphism group intersects a basic open set $N_{\bar{a},\bar{b}}$ if and only if the substructures $\bar{a}$ and $\bar{b}$ are isomorphic. Thus the map from the Fraïssé structures to their automorphism groups is Borel. ∎
Corollary 6.3.
There are continuum many pairwise non-isomorphic Polish groups that are extremely amenable. There are continuum many pairwise non-isomorphism separable metrizable groups what are not Polish and are extremely amenable.
Proof.
For both results, note that
$_{\mathbb{R}}\ <_{B}\,E_{0}\ \leq_{B}E^{IR}_{GL_{2}(\mathbb{Q})}\ \leq_{B}\ \approx_{\mathsf{CDVS}_{0}}$
by Lemma 5.11 and Theorem 5.10. Thus there are continuum many pairwise $\approx_{\mathsf{CDVS}_{0}}$-inequivalent elements. Now the first conclusion follows from Theorem 4.3, and the second conclusion follows from Theorems 4.2 and 3.4 (1). ∎
We mention some open problems.
Let us first mention another class of automorphism groups known to be extremely amenable. Recall that an ultrametric is a metric satisfying $d(x,z)\leq\max\{d(x,y),d(y,z)\}$ for any $x,y,z$. An ordering << on an ultrametric space $X$ is convex if for any $x\in X$ and $s>0$, the ball $B(x,s)=\{y\in X:d(x,y)\leq s\}$ is a convex set in the ordering <<, i.e. if $y_{1}<z<y_{2}$ and $y_{1},y_{2}\in B(x,s)$, then $z\in B(x,s)$. An ultrametric space with a convex ordering is called a convexly-ordered ultrametric space.
@@
Denoting the Fraïssé limit of this class by $B_{S}^{c<}$, by the KPT correspondence, ${\rm Aut}(B_{S}^{c<})$ is extremely amenable. We note that if $S$ and $T$ are countable sets of positive real numbers that are order-isomorphic, then ${\rm Aut}(B_{S}^{c<})$ and ${\rm Aut}(B_{T}^{c<})$ are isomorphic as topological groups.
Question 6.5.
Suppose $S$ and $T$ are countable sets of positive real numbers such that ${\rm Aut}(B_{S}^{c<})$ and ${\rm Aut}(B_{T}^{c<})$ are isomorphic as topological groups. Are $S$ and $T$ order-isomorphic?
If the answer is positive, then we would obtain another proof of Theorem 6.2.
We note that there is a prominent difference between Theorems 1.2 and 1.5, namely, in Theorem 1.2 we have the equivalence between the isomorphism of the topological groups and the isomorphism of them as abstract groups, while in Theorem 1.5 we do not have this equivalence. The reason for this equivalence to hold in Theorem 1.2 is that $\mbox{\rm Aut}(\mathbb{U}_{\Delta})$ has ample generics (see Corollary 4.1 of Solecki [34]), i.e., $G$ has a comeager orbit in the conjugacy action of $G$ on $G^{n}$, for each $n\geq 1$. By a result of Kechris–Rosendal [18], if a Polish group has ample generics, then every abstract homomorphism into a separable group is automatically continuous and thus it has a unique Polish topology. However, for the ordered case, Slutsky [33] showed that ${\rm Aut}(\mathbb{U}_{\Delta}^{<})$ does not have ample generics. Hence the following question remains.
Question 6.6.
[…]
Of course, the main question left open by this paper is the following.
Question 6.7.
What is the exact complexity of $\cong_{\mathsf{EA}}$ in the Borel reducibility hierarchy?
Recall from the Introduction that there are many examples of extremely amenable Polish groups of the form $L_{0}(\mu,G)$, and in particular many of them are abelian. However, there is currently no known method to distinguish their isomorphism types. Thus we still have the following question.
Question 6.8.
Are threre continnum many pairwise non-isomorphic extremely amenable abelian Polish groups?